From nch@roe.ac.uk Fri Feb 19 17:18:08 2010 Date: Fri, 19 Feb 2010 12:36:33 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS science archives weekly project meeting minutes, 19/2/10 Minutes of WFAU VDFS Science Archive meeting: February 19th 2009 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, PMW, ETWS, NJC, RGM, RSC, RPB, MSH, MAR Apologies: JPE, AL, KTN DoNM: 10am, Friday February 26th 2010 in the VISTA Hut Actions discharged this week: ----------------------------- ACTION: MAR, ETWS & NJC to put summary notes of the PSPIs meeting on the TWiki. Discharged ACTION: NCH and RSC to meet next week to sort out C++ implementation of proper motions in the merged source lists for UKIDSS Discharged Actions partly discharged but continuing: ----------------------------------------- ACTION: RPB & NJC (with help when required from RSC) set up cross-neighbour tables between Stripe 82 and WSA/VSA surveys as appropriate. Done for the WSA (LAS & DXS have coverage in common); it was noted that the same should be done for VISTA VHS, VIKING (& possibly VIDEO) in the VSA. Actions carried forward from 12/02/10 meeting: ---------------------------------------------- None this week Specific points and new actions: -------------------------------- Project management: The team agreed to meet again on Wed 24th at 2pm in the Vista Hut to brain-storm the necessary operational and SW changes falling requests as a result of the PSPIs meeting. ACTION: NCH to book the Vista Hut meeting room for 2pm Wed 24th. WFCAM & VISTA updates: Nothing new to report this week. Comments and issues arising from CASU minutes: None this week. Networking: Some JANET maintenance messages over the last week, but otherwise all quiet on the network front. WSA/VSA Operations: RPB reported that UKIDSS DR7+ is ready to go - MAR and NCH volunteered to do some final QC2/sanity checks. (SJW has checked the areal coverage of the LAS and GCS from the source merging logs.) ETWS noted: - Re-created broken JPGs for 05A to 08B. - Created browser pages for latest non-survey release. - Created metadata docs of SDSS BestDR5 and BestDR7. - More work on automation of JPG quality control. - Tested multiple fits files for JRL to locate problem in FITS key updates of VISTA data, but no difference in timing was found between C/cfitsio code and pyfits. Hardware and Systems: MSH noted that hatshepsut will be moved into the cluster rack today and will become the new SQL2008 test DB server configured a la ramses operational nodes. RSC asked that G:H:I: partitions be set up for equivalence with these. Invitations to tender for a new curation server have been invited; an LTO4 drive for the tape library unit has been ordered. Software: NJC reported: "This week I have finished testing the new template schema for VISTA surveys and non-surveys that will allow the surveys to be run through ProgrammeBuilder. I still have to tidy up ProgrammeBuilder and write some notes on the Wiki page." MAR has fettled the VSA User Interface for a few changes requested for the VMC; some issues have arisen over truncated strings read from FITS files so ETWS will be reviewing SQL VARCHAR types in the current VSA schema definition. RSC noted * CU19 has been updated to include calibration metadata in VISTA release databases. * I've started implementing the proper motion calculations in CU7. Survey Data Release: As noted above, UKIDSS DR7+ is ready for release DB creation; final checks to be finished by close of play today then CU19 will be set running overnight. Non-survey Data Release: Further 09A releases have taken place this week, including a Korean project. Requests for 09B products in advance of routine operations are also being looked after on a best-efforts basis. Astrogrid deployment & Data Analysis services: MSH noted that the venerable VO server is rapidly becoming overloaded what with mutiple applications services. The team agreed that it is important to have a beefy unit for WSA/VSA VO services, so a replacement will be ordered. All VO services have had updated metadocs created and are now exposed through the new TAP service; secure conesearch for UKIDSS DR4+ et seq have been set up (although there is a dearth of suitable client-end software). RSC has looked into the list-driven photometry tool (co-incidentally some new problems have been reported to wsa-support today). Issues with results emails going missing may be due to recent reorganisation of mail services on site, but investigations continue (necessarily at a low priority unfortunately...) Miscellaneous: It was agreed that WFAU should have strong representation at the Glasgow NAM: http://www.astro.gla.ac.uk/nam2010/ ACTION: All the consider possible poster/talk contributions by next Friday at which point we'll have a look-see and plug any gaps. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================